Are the following vectors linearly independent?

  1. (1, 2), (-2, 4)
  2. (0, 0, 1), (0, 1, 1), (1, 1, 1)
  3. (1, 2), (1, 3), (1, 1)
  4. (1, 2, 3), (2, 3, 4)
  5. (0, 1, 2), (2, 0, 1), (0, 0, 1), (3, 2, 1)
1
Expert's answer
2021-11-30T11:23:54-0500


1. Since "\frac{1}{2}\ne\frac{-2}{4}," the vectors "(1, 2), (-2, 4)" are not collinear, and thus they are linearly independent.


2. Consider the linear combination "a(0, 0, 1)+b (0, 1, 1)+c (1, 1, 1)=(0,0,0)," which is equivalent to "(c,b+c,a+b+c)=(0,0,0)," and hence "a=b=c=0." We conclude that the vectors "(0, 0, 1), (0, 1, 1), (1, 1, 1)" are linear independent.


3. Since in 2-dimentinal space any three vectors are linearly dependent, we conclude that the vectors "(1, 2), (1, 3), (1, 1)" are linearly dependent.


4. Since "\frac{1}{2}\ne\frac{2}{3}\ne\frac{3}4," we conclude that the vectors"(1, 2, 3), (2, 3, 4)" are not collinear, and hence they are linearly independent.


5. Since in 3-dimentinal space any four vectors are linearly dependent, we conclude that the vectors "(0, 1, 2), (2, 0, 1), (0, 0, 1), (3, 2, 1)" are linearly dependent.